# Benchmarking Time Series Foundation Models for Short-Term Household Electricity Load Forecasting

### v2

@@ 3.2 Exploratory Data Analysis (EDA)
Figure 3 shows the autocorrelation for a single house’s median electricity load per hour in Southern Germany during March 2016. A cyclical pattern is visible in the time series in a sinuous form, also described in Mouakher et al. (2022) for 24-hour lags, where the energy consumption for an hour is strongly correlated with the same hour of the following day. Daily patterns and weekly patterns are common for many houses in all datasets.
@@
The datasets have individual missing values between two measurements or, in the case of the Lower Saxony and REFIT datasets, up to weeks-long gaps. The reasons given in the datasets’ descriptions are technical failures of the data logger, Internet outages, conversion work Schlemminger et al. (2022), failed radio transmissions, and problems in daylight savings time transitions Open Power System Data (2020).
@@
3.3 Data Preprocessing
3.3.1 Handling Missing Values
@@
3.3.2 Train-Test Split
@@
Using the train-test method proposed by Hertel et al. (2023), splitting the data into 70 % training, 10 % validation, and 20 % testing would result in different split dates for each household in our datasets. This presents a risk of information leakage, for instance, global time-specific patterns, such as the Covid-19 crisis, could be represented in one time series training data (Household A). Another time series (Household B) could have a different split date with a test set that comprises the same time period as the training data of other households (Household A). When using this approach, the global pattern might be learned (training data Household A) by global models and transferred to other time series (Household B). To safeguard against the leakage of information, we select a unique split-date for each dataset. Additionally, we ensure no overlapping of the datasets by cutting them if the evaluation time frame of one dataset overlaps with the training data of another dataset.
@@
Figure 6: Train-Test split. Time spans of each time series
First, we determine the time point $t_{0.25}$ of the 0.25th percentile of each household’s maximum date (Step 1). In Step 2, we calculate all possible time points $T_{poss}=(t_{min},...,t_{0.25})$ (i.e., hours) between the global minimum date $t_{min}$ and $t_{0.25}$, regardless of their frequency in the actual data. Based on this, we determine the final train-test split at the 0.8th percentile of $T_{poss}$ (Step 3). By ensuring that the date defined in Step 1 is included in our test set, we ensure that the test set comprises a large number of households.
@@
3.4 Evaluated Models
A condition for the selection of TSFM was the disclosure of the training data and the provision of a pre-trained model as discussed in chapter 3.1. At the time of writing, Chronos Ansari et al. (2024), TimesFM Das et al. (2024), Sundial Liu et al. (2025), Time-MoE Shi et al. (2024), Moirai Woo et al. (2024) and LagLlama Rasul et al. (2024) fulfill both conditions. The TSFM are not be fine-tuned on the household datasets introduced for our evaluation. Only inference is used on the unseen data for achieving zero-shot predictions Xian et al. (2019). Therefore we excluded models which need finetuning e.g. the TSFM Moment Goswami et al. (2024), which forecasting head is randomly initialized. The implemention is done by following the suggested usage from the official Github repository and using the provided model weights. For LagLlama, we activate RoPe scaling as suggested in the zero-shot tutorial for input sizes longer than the context length of 32, because most of the tested input sizes are above that threshold. As the main TFS competitors for the foundation models, we used time series models with a similar Transformer-based architecture. Namely, we chose the original Transformer Encoder-Decoder implementation ("VanillaTransformer") Vaswani et al. (2017), and PatchTST Nie et al. (2023), which represents the current SOTA for household energy STLF (see Table 1). Furthermore, we chose iTransfomer Liu et al. (2024) and Temporal Fusion Transformer (TFT) Lim et al. (2021), which we identified in our literature review. The TFS Transformers are trained univariate, which can be a limiting factor for the TFS Transformers to varying degrees. However, this approach allows for the most direct comparison with the foundation models. For implementing the TFS Transformers, we used the NeuralForecast library, which is referring its model implementations to the corresponding papers.  [^1] The complete preprocessing, training and evaluation code can be found under this anonymous github repo: https://anonymous.4open.science/r/benchmarking_tsfm_household_load_forecasting-DF75/
As a baseline, we take the SeasonalAverage Hyndman and Athanasopoulos (2018) with a seasonality of 24 as we observed a high autocorrelation in the data with lags of 24. The SeasonalAverage is calculated per time series with the Pandas library with an hourly average over the input size as forecasts for the horizon. Additionally, we report as baseline reference the Naive Forecast where the last observed value is the forecast for the complete forecast horizon.
@@ 3.4 Evaluated Models
A complete overview of the used models and hyperlinks to their implementation can be found in Table 3.4.
Table 5: Comparison of models, their architectures and implementations
@@
3.5 Training and Evaluation
We follow the approach of a time series cross-validation Bergmeir and Benítez (2012) with a calibration window. The TFS Transformers are trained on a fixed window size (365 days) of the most recent observations up to the split date. After the initial forecast, both the split date and the calibration window are moved forward by the duration of the horizon. Consequently, the training data now encompasses the most recent 365 days, starting from this updated split date. The TFS models are retrained from scratch on the updated calibration window, generating new forecasts for the next horizon period.
@@ 3.5 Training and Evaluation
Following the insights that the global is the best training approach Hertel et al. (2023), we train the TFS Transformers on all the datasets’ training splits together.
@@
3.6 Challenges in Evaluating Foundation Models
While we prevent information leakage of global temporal patterns in the globally trained TFS Transformer models, there remains a special challenge in evaluating time series foundation models like Chronos Ansari et al. (2024), TimesFMDas et al. (2024) or LagLlama Rasul et al. (2024). The pre-trained models might have learned global temporal patterns (e.g., Covid-19 Prabowo et al. (2023), geopolitical crises) due to the massive amount of data used during training. Any evaluation on a dataset that is in the same date range as the training data of the foundation model could be potentially affected by such information leakage. Even cross-domain influences are possible, e.g., as weather affects household energy consumption Kang and Reiner (2022), a foundation model trained on historical weather data could theoretically lead to information leakage on a household energy dataset in the same time period.
@@ 3.7 Metrics
  • $\hat{y}_{h,i}$ is the prediction for household $h$ for prediction $i$ out of $n$
  • $y_{h,i}$ is the actual value for household $h$ for prediction $i$ of $n$
  • $h$ is the number of households
  • $n$ is the number of predictions per household
The same logic is applied to compute the Mean Squared Error per household (MSEh). In addition, we use the adjusted p-norm error per house (APNEh), introduced by Haben et al. (2014), which is specifically designed to address the “double penalty” effect. This effect occurs when a forecast that correctly predicts the magnitude of the target (such as a peak) but is slightly displaced in time and is penalized more heavily than a constant, less informative forecast. Traditional point-wise metrics like MAE or MSE fail to account for such temporal misalignments. The adjusted p-norm error mitigates this by searching for a restricted temporal permutation of the forecast that minimizes the error according to a specified p-norm. This search is constrained by an adjustment limit window w, which defines the maximum allowable shift between the forecasted and actual time points. Following the recommendation of the authors, we use a p-norm of 4 and only adapt the adjustment window from w = 3 to w = 1 as we have hourly data instead of half-hourly data.
Based on the various occurrences of zero or near-zero values, we discard other metrics, such as the mean percentage error (MAPE) and symmetric MAPE (SMAPE), as they tend to increase drastically with values near to zero Hyndman and Koehler (2006).
@@
4 Results
Table 6: Model’s MAEh, MSEh and APNEh scores across all datasets
@@ 4 Results
Table 6 shows the MAEh, MSEh and APNEh for all models across the different datasets. Comparing the different metrics, there is no single model which dominates the benchmark. But stepping back, we can see that the familiy of TSFM is outperforming the other approaches in most cases. Except for the IDEAL dataset, all best or second best results are achieved by TSFM. For the MAEh Chronos-Bolt delivers best performances across the Lower Saxony, REFIT and Southern Germany datasets, often with TimesFM following closely. The newer version TimesFM 2.0 outperformed Chronos-Bolt on the IDEAL dataset. But also the TFS Transformer PatchTST is performing good or even outperforming the other models on the IDEAL dataset in terms of MSEh. On the other datasets Time-MoE and Sundial have the best performance for the MSEh metric. The worst-performing models overall were Lag-LLama and Moirai, which could not outperform the SeasonalAverage baseline.
Table 7: Rank (MAEh) and MAEh results for every input size and horizon for all datasets. Best results are in bold, second best underlined.
@@
A more detailed analysis how the models behave with different input sizes and horizons allows Table 7, which shows the average rank based on MAEh and the MAEh by model, input size, and horizon.
5 Discussion
@@
LagLlama performs poorly for all input sizes compared to the other foundation models and most TFS Transformers. This might be explained by LagLlama’s architecture, which uses lags to predict future values Rasul et al. (2024). These lags include quarterly, monthly, weekly, daily, and hourly levels Rasul et al. (2024), while our defined maximum input size of 168 hours allows only incorporating daily to weekly lags. Similarly our restricted experimental setup with limited information about the time series could also explain the performance of the Moirai model, which was originally evaluated with a significantly longer context size of 1000 Woo et al. (2024).
Considering the performance of the TFS Transformer PatchTST, our study supports the findings of Hertel et al. (2023) and Cen and Lim (2024), as the PatchTST model provides good overall performance and even slightly outperforms other TSFM on an input size of 24 hours. Additionally, the results of Hertel et al. (2023) indicate that TFS Transformers perform better when trained globally with more data. Our study extends their findings by showing that the zero-shot performance of the also globally trained TSFM is better or comparable to TFS Transformers. Interestingly, we could not reproduce the statement that input sizes exceeding the horizon decreases forecasting performance of transformer models Wen et al. (2022), but observed the opposite: PatchTST as well as all TSFM benefit from longer context. Just the the VanillaTransformer did not benefit from longer input sizes matching the results of Wen et al. (2022).
@@
Compared to standard TSFM evaluations (see Section 2), our approach combines multi-dataset evaluation with time-series cross-validation for a more comprehensive assessment. This design mitigates the limitations of smaller datasets by capturing both cross-series variation and performance changes over time, thereby strengthening the validity of our findings and supporting stronger claims about the robustness and generalizability of TSFMs relative to the TFS Transformer.
6 Limitations & Outlook
@@
Naturally, our analysis is not without limitations. Hyperparameter tuning the TFS Transformers may increase the performance of these models supporting the findings of Sievers and Blank (2023) and Upadhyay et al. (2023) while simultaneously increasing the training effort compared to TSFM even more. On the other hand, longer input sizes could lead to better performance of the foundation models, especially for LagLlama and Moirai, which might need higher context lengths due to their models architecture Rasul et al. (2024); Woo et al. (2024).
@@
There are several directions for future research on household STLF. First, incorporating longer input sizes would allow for drawing a better picture of TFS Transformer-based and foundation model behavior, e.g., the behavior of LagLlama on longer inputs. While TSFM provide zero-shot forecasts that can outperform SOTA TFS Transformers, fine-tuning foundation models has been shown to increase performance in other disciplines, such as foundation LLMs. Moreover TSFM, which have been trained on some household energy data, performed better than other TSFMs. Hence, pre-training and fine-tuning foundation models on energy time series forecasting could further increase their capabilities. Second, we propose that the models can learn cross-domain global patterns, such as the COVID-19 pandemic or geopolitical crises, which may lead to information leakage when evaluating these models on holdout datasets originating from the same time period as the training datasets. This possible problem should be explored in future research. Third, our univariate analysis could be extended to multivariate TSFMs, including covariates such as weather data Kang and Reiner (2022).
7 Conclusion
@@
Motivated by recent algorithmic developments in time series forecasting, this study investigated whether time series foundation models are competitive to SOTA TFS Transformers on household STLF tasks. Following the guideline of Haben et al. (2021) we considered time series cross-validation on multiple datasets using statistical baselines and sophisticated TFS Transformers. In our benchmark, unlike the TFS Transformers, the TSFMs were used out-of-the-box for prediction without task-specific adaptation or fine-tuning. We show that TSFM are already capable of delivering competitive and in most cases better forecast performance compared to trained-from-scratch Transformers. The foundation models show their strength, especially when there is more context (i.e., longer input sizes). Furthermore, our findings indicate that, in the case of LagLLama or Moirai, its special architecture may harm performance when dealing with a limited input context. In conclusion, the ability of foundation models to achieve high accuracy with limited data and without training opens up new possibilities for developing more efficient and accessible energy forecasting solutions.

### v3

@@ III-B Exploratory Data Analysis (EDA)
Figure 3 shows the autocorrelation for a single house’s median electricity load per hour in Southern Germany during March 2016. A cyclical pattern is visible in the time series in a sinuous form, also described in [Mouakher2022] for 24-hour lags, where the energy consumption for an hour is strongly correlated with the same hour of the following day. Daily patterns and weekly patterns are common for many houses in all datasets.
@@
The datasets have individual missing values between two measurements or, in the case of the Lower Saxony and REFIT datasets, up to weeks-long gaps. The reasons given in the datasets’ descriptions are technical failures of the data logger, Internet outages, conversion work [Schlemminger2022], failed radio transmissions, and problems in daylight savings time transitions [OPSD_2020].
@@
III-C Data Preprocessing
III-C1 Handling Missing Values
@@
III-C2 Train-Test Split
Figure 6: Train-Test split. Time spans of each time series
@@ III-C2 Train-Test Split
Using the train-test method proposed by Hertel.2023, splitting the data into 70 % training, 10 % validation, and 20 % testing would result in different split dates for each household in our datasets. This presents a risk of information leakage, for instance, global time-specific patterns, such as the Covid-19 crisis, could be represented in one time series training data (Household A). Another time series (Household B) could have a different split date with a test set that comprises the same time period as the training data of other households (Household A). When using this approach, the global pattern might be learned (training data Household A) by global models and transferred to other time series (Household B). To safeguard against the leakage of information, we select a unique split-date for each dataset. Additionally, we ensure no overlapping of the datasets by cutting them if the evaluation time frame of one dataset overlaps with the training data of another dataset.
@@
[…]
@@
III-D Evaluated Models
A condition for the selection of TSFM was the disclosure of the training data and the provision of a pre-trained model as discussed in chapter III-A. At the time of writing, Chronos [Ansari.2024], TimesFM [Das2024], Sundial [liu2025sundial], Time-MoE [shi_time-moe_2024], Moirai [woo_unified_2024] and LagLlama [Rasul.2024] fulfill both conditions. The TSFM are not be fine-tuned on the household datasets introduced for our evaluation. Only inference is used on the unseen data for achieving zero-shot predictions [zero-shot-xian2019]. Therefore we excluded models which need finetuning e.g. the TSFM Moment goswami_moment_2024, which forecasting head is randomly initialized. The implemention is done by following the suggested usage from the official Github repository and using the provided model weights. For LagLlama, we activate RoPe scaling as suggested in the zero-shot tutorial for input sizes longer than the context length of 32, because most of the tested input sizes are above that threshold. As the main TFS competitors for the foundation models, we used time series models with a similar Transformer-based architecture. Namely, we chose the original Transformer Encoder-Decoder implementation ("VanillaTransformer") [Vaswani.2017], and PatchTST [nie2023time], which represents the current SOTA for household energy STLF (see Table I). Furthermore, we chose iTransfomer [Liu.2024] and Temporal Fusion Transformer (TFT) [lim2021temporal], which we identified in our literature review. The TFS Transformers are trained univariate, which can be a limiting factor for the TFS Transformers to varying degrees. However, this approach allows for the most direct comparison with the foundation models. For implementing the TFS Transformers, we used the NeuralForecast library, which is referring its model implementations to the corresponding papers.  [^1] The complete preprocessing, training and evaluation code can be found under this github repo: https://github.com/mmcux/benchmarking_tsfm_household_load_forecasting
As a baseline, we take the SeasonalAverage [hyndman2018forecasting] with a seasonality of 24 as we observed a high autocorrelation in the data with lags of 24. The SeasonalAverage is calculated per time series with the Pandas library with an hourly average over the input size as forecasts for the horizon. Additionally, we report as baseline reference the Naive Forecast where the last observed value is the forecast for the complete forecast horizon.
@@ III-D Evaluated Models
A complete overview of the used models and hyperlinks to their implementation can be found in Table III-D.
TABLE V: Comparison of models, their architectures and implementations
@@
III-E Training and Evaluation
We follow the approach of a time series cross-validation [Bergmeir.2012cross-validation] with a calibration window. The TFS Transformers are trained on a fixed window size (365 days) of the most recent observations up to the split date. After the initial forecast, both the split date and the calibration window are moved forward by the duration of the horizon. Consequently, the training data now encompasses the most recent 365 days, starting from this updated split date. The TFS models are retrained from scratch on the updated calibration window, generating new forecasts for the next horizon period.
@@ III-E Training and Evaluation
Following the insights that the global is the best training approach [Hertel.2023], we train the TFS Transformers on all the datasets’ training splits together.
@@
III-F Challenges in Evaluating Foundation Models
While we prevent information leakage of global temporal patterns in the globally trained TFS Transformer models, there remains a special challenge in evaluating time series foundation models like Chronos [Ansari.2024], TimesFM[Das2024] or LagLlama [Rasul.2024]. The pre-trained models might have learned global temporal patterns (e.g., Covid-19 [Prabowo2023], geopolitical crises) due to the massive amount of data used during training. Any evaluation on a dataset that is in the same date range as the training data of the foundation model could be potentially affected by such information leakage. Even cross-domain influences are possible, e.g., as weather affects household energy consumption [kang2022weatherelectricityconsumption], a foundation model trained on historical weather data could theoretically lead to information leakage on a household energy dataset in the same time period.
@@ III-G Metrics
[…]
  • hh is the number of households
  • nn is the number of predictions per household
The same logic is applied to compute the Mean Squared Error per household (MSEh). In addition, we use the adjusted p-norm error per house (APNEh), introduced by haben2014new, which is specifically designed to address the “double penalty” effect. This effect occurs when a forecast that correctly predicts the magnitude of the target (such as a peak) but is slightly displaced in time and is penalized more heavily than a constant, less informative forecast. Traditional point-wise metrics like MAE or MSE fail to account for such temporal misalignments. The adjusted p-norm error mitigates this by searching for a restricted temporal permutation of the forecast that minimizes the error according to a specified p-norm. This search is constrained by an adjustment limit window w, which defines the maximum allowable shift between the forecasted and actual time points. Following the recommendation of the authors, we use a p-norm of 4 and only adapt the adjustment window from w = 3 to w = 1 as we have hourly data instead of half-hourly data.
Based on the various occurrences of zero or near-zero values, we discard other metrics, such as the mean percentage error (MAPE) and symmetric MAPE (SMAPE), as they tend to increase drastically with values near to zero [Hyndman2006].
[…]
@@
TABLE VI: Model’s MAEh, MSEh and APNEh scores across all datasets
@@
IV Results
[…]
For the MAEh Chronos-Bolt delivers best performances across the Lower Saxony, REFIT and Southern Germany datasets, often with TimesFM following closely. The newer version TimesFM 2.0 outperformed Chronos-Bolt on the IDEAL dataset. But also the TFS Transformer PatchTST is performing good or even outperforming the other models on the IDEAL dataset in terms of MSEh. On the other datasets Time-MoE and Sundial have the best performance for the MSEh metric. The worst-performing models overall were Lag-LLama and Moirai, which could not outperform the SeasonalAverage baseline.
A more detailed analysis how the models behave with different input sizes and horizons allows Table VII, which shows the average rank based on MAEh and the MAEh by model, input size, and horizon.
TABLE VII: Rank (MAEh) and MAEh results for every input size and horizon for all datasets. Best results are in bold, second best underlined.
@@ IV Results
V Discussion
@@
LagLlama performs poorly for all input sizes compared to the other foundation models and most TFS Transformers. This might be explained by LagLlama’s architecture, which uses lags to predict future values [Rasul.2024]. These lags include quarterly, monthly, weekly, daily, and hourly levels [Rasul.2024], while our defined maximum input size of 168 hours allows only incorporating daily to weekly lags. Similarly our restricted experimental setup with limited information about the time series could also explain the performance of the Moirai model, which was originally evaluated with a significantly longer context size of 1000 [woo_unified_2024].
Considering the performance of the TFS Transformer PatchTST, our study supports the findings of Hertel.2023 and Cen.2024, as the PatchTST model provides good overall performance and even slightly outperforms other TSFM on an input size of 24 hours. Additionally, the results of Hertel.2023 indicate that TFS Transformers perform better when trained globally with more data. Our study extends their findings by showing that the zero-shot performance of the also globally trained TSFM is better or comparable to TFS Transformers. Interestingly, we could not reproduce the statement that input sizes exceeding the horizon decreases forecasting performance of transformer models [wen_Transformers_survery_2022], but observed the opposite: PatchTST as well as all TSFM benefit from longer context. Just the the VanillaTransformer did not benefit from longer input sizes matching the results of [wen_Transformers_survery_2022].
@@
Compared to standard TSFM evaluations (see Section II), our approach combines multi-dataset evaluation with time-series cross-validation for a more comprehensive assessment. This design mitigates the limitations of smaller datasets by capturing both cross-series variation and performance changes over time, thereby strengthening the validity of our findings and supporting stronger claims about the robustness and generalizability of TSFMs relative to the TFS Transformer.
VI Limitations & Outlook
@@
Naturally, our analysis is not without limitations. Hyperparameter tuning the TFS Transformers may increase the performance of these models supporting the findings of Sievers.2023 and Upadhyay.2023 while simultaneously increasing the training effort compared to TSFM even more. On the other hand, longer input sizes could lead to better performance of the foundation models, especially for LagLlama and Moirai, which might need higher context lengths due to their models architecture [Rasul.2024, woo_unified_2024].
@@
There are several directions for future research on household STLF. First, incorporating longer input sizes would allow for drawing a better picture of TFS Transformer-based and foundation model behavior, e.g., the behavior of LagLlama on longer inputs. While TSFM provide zero-shot forecasts that can outperform SOTA TFS Transformers, fine-tuning foundation models has been shown to increase performance in other disciplines, such as foundation LLMs. Moreover TSFM, which have been trained on some household energy data, performed better than other TSFMs. Hence, pre-training and fine-tuning foundation models on energy time series forecasting could further increase their capabilities. Second, we propose that the models can learn cross-domain global patterns, such as the COVID-19 pandemic or geopolitical crises, which may lead to information leakage when evaluating these models on holdout datasets originating from the same time period as the training datasets. This possible problem should be explored in future research. Third, our univariate analysis could be extended to multivariate TSFMs, including covariates such as weather data [kang2022weatherelectricityconsumption].
VII Conclusion
@@
Motivated by recent algorithmic developments in time series forecasting, this study investigated whether time series foundation models are competitive to SOTA TFS Transformers on household STLF tasks. Following the guideline of Haben.2021 we considered time series cross-validation on multiple datasets using statistical baselines and sophisticated TFS Transformers. In our benchmark, unlike the TFS Transformers, the TSFMs were used out-of-the-box for prediction without task-specific adaptation or fine-tuning. We show that TSFM are already capable of delivering competitive and in most cases better forecast performance compared to trained-from-scratch Transformers. The foundation models show their strength, especially when there is more context (i.e., longer input sizes). Furthermore, our findings indicate that, in the case of LagLLama or Moirai, its special architecture may harm performance when dealing with a limited input context. In conclusion, the ability of foundation models to achieve high accuracy with limited data and without training opens up new possibilities for developing more efficient and accessible energy forecasting solutions.